OpenStudy (anonymous):

Can you show me step by step negative 8 plus the sqare root of 32 over 24

OpenStudy (unklerhaukus):

do you mean\[-8+\sqrt{\frac{32}{24}}\] or \[-8+{\frac{\sqrt{32}}{24}}\] or \[\frac{-8+\sqrt{32}} {24}\]